OpenStudy (anonymous):

Solve for m. 2m + 3 = 7 m = 5 m = 4 m = 10.5 m = 2

pooja195 (pooja195):

whats 7-3?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

subtract 3 from 7 then divide it by 2

OpenStudy (anonymous):

It's D.. Cause its two

OpenStudy (anonymous):

u sure

OpenStudy (anonymous):

Haha... Not really.

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

2m + 3 = 7 --- subtract 3 from both sides 2m + 3 - 3 = 7 - 3 2m = 4 -- now divide both sides by 2 m = 4/2 m = 2 you can even check your answer 2(2) + 3 = 7 4 + 3 = 7 7 = 7 (correct) so m = 2

OpenStudy (anonymous):

Okay, so I did do the math correct... Thank you :)